OpenStudy (kaleidoscopicsink):

Implicit derivative of ysin(x) = xsin(y). I know what the answer is: (ycos(x)-sin(y))/ (xcos(y) - sin(x)) Can someone just show the steps for the implicit differentiation?

OpenStudy (kenshin):

sinxdy/dx + ycosx = siny + xcosy dy/dx; siny-ycosx = dy/dxsinx - xcosydy/dx; factor out dy/dx and divide xsinx - xcosy from both sides

OpenStudy (kaleidoscopicsink):

I need step by step. Like shown to me.

OpenStudy (kaleidoscopicsink):

Seeing that does not really help my understanding.

OpenStudy (kenshin):

product rule between y and sinx, product rule between x and sin(y); when doing derivative of sinx and x you get their derivative straight away; when doing derivative of y and sin(y) you gotta times it by dy/dx coz it's implicit diff

OpenStudy (kaleidoscopicsink):

I need to see it step by step like each individual one wrote out. Like 2x+2=4 2x=4-2 --> 2x=2 --> x=2/2--> x=1. Words mean nothing, I need the actual math part elaborated to understand.

OpenStudy (kenshin):

dy/dy * dy/dx = dy/dx; d(siny)/dy * dy/dx = d(siny)/dx, those are the concepts around why you have to times by dy/dx by the derivative of siny and y with respect to y.

OpenStudy (irishboy123):

so we start with: \(\dfrac{d}{dx} ysin(x) \)

OpenStudy (kaleidoscopicsink):

I do not understand that. Can you show me the math, please? I'm just getting more confused.

OpenStudy (irishboy123):

K sorry i can't be of more help. but @ikram002p is really good at explaining stuff, ..... and she's on the thread!!!

OpenStudy (ikram002p):

ysin(x) = xsin(y). @KaleidoscopicsInk start with the right side, could you derive it with respect to x ?

OpenStudy (kaleidoscopicsink):

I have no idea how to.

OpenStudy (ikram002p):

(x sin(y))' fisrt u can use this hint (ab)'=a'b+ab'

OpenStudy (kaleidoscopicsink):

In all honesty, I just need someone to show me this problem step by step. Im tired and just need this last one done. I have no means of retaining the information. I just want this to be over with. I can do derivatives but I lack the patience to do them for a class that means absolutely nothing to me because it is honestly one of the most retarded and slow classes I have been in. If you could just do me one favor, out of the kindness of your heart and just show me how to do this problem, i'd be forever grateful.

OpenStudy (ikram002p):

so (x sin y)'=x'(sin y)+x (sin y') x'=1 (sin y)' you can do this by the implicit differentiation rule which is [f(g(x))]'=f' * g' thus (sin y)'=cos y* y' y'=dy/dx right ?

OpenStudy (kaleidoscopicsink):

Sure. I really don't know. I just want this over with, please.

OpenStudy (ikram002p):

well, i don't think your interesting in learning this. and i'm sure i'm not interested in teaching this to people who keep whining and not listening. i'm trying to do step by step solution to u and u are like (please i wanna this over). hold your self up and try to focus WOULD YOU!

OpenStudy (kaleidoscopicsink):

As hard as it is to believe, I am. I have been. I'm trying not to break at the moment, so sittting here and being confused and nearly lectured is difficult to bare because all I want is this class to be over and my head is killing me. I'm driving myself insane with this one question. I'm trying, if you believe it or not, that's your choice. But I'm freaking out and my calm is completely gone. I'm sorry. I am if I come off that way. Im not trying to be rude, I'm trying not to break.

OpenStudy (ikram002p):

just keep calm and take a rest =). it doesn't worth all that stress for real. back to the topic now " implicit differentiation " is used when you two or more variables sometimes they are not separable.

OpenStudy (kaleidoscopicsink):

I know, I can't rest. I get yelled at if I do.

OpenStudy (kaleidoscopicsink):

I know what implicit differentiation is.

OpenStudy (ikram002p):

you know what it is. good ! what is the derivative of sin(x) ?

OpenStudy (kaleidoscopicsink):

cos(x)

OpenStudy (ikram002p):

good! and what is derivative of sin(y) with respect to x (dy/dx) ?

OpenStudy (kaleidoscopicsink):

I don't know.

OpenStudy (ikram002p):

use this rule f(y)'=f'(y)*y'

OpenStudy (ikram002p):

give a try if you couldn't find it i'll tell you how.

OpenStudy (kaleidoscopicsink):

cos(x)sin(y)

OpenStudy (ikram002p):

very close but y'=dy/dx

OpenStudy (kaleidoscopicsink):

I don't know.

OpenStudy (ikram002p):

well sin(y)'=cos y*y'=cos y *(dy/dx)

OpenStudy (kaleidoscopicsink):

And?

OpenStudy (ikram002p):

\(\Large {\color{Red} { \text{remember } }} (ab)'=ab'+a'b \\ \begin{matrix} (y~\sin(x))' = (x ~\sin(y) )' \\ y(\sin (x))'+y'\sin(x)=x(\sin(y))'+x'(\sin(y)) \\ y \cos(x)+\dfrac{dy}{dx} \sin (x) =x \cos(y)\dfrac{dy}{dx} + \sin (y) \\ \end{matrix} \)

OpenStudy (kaleidoscopicsink):

Okay, I know the product rule.

OpenStudy (ikram002p):

seems u know everything now cool !

OpenStudy (kaleidoscopicsink):

Can you show me how to do the problem now?

OpenStudy (ikram002p):

well check the last line we have concluded lately \( \large y \cos(x)+{\color{Red} {\dfrac{dy}{dx}}} \sin (x) =x \cos(y){\color{Red} {\dfrac{dy}{dx}}} + \sin (y) \)

OpenStudy (ikram002p):

all you need now is to put dy/dx on a side on it's own could you do that ?

OpenStudy (kaleidoscopicsink):

No.

OpenStudy (ikram002p):

assume you have this problem \(\Large a+ b\times c = d\times c +e\) can you solve for c ?(put c on a side on it's own?)

OpenStudy (kaleidoscopicsink):

I cannot. I'm getting yelled at. My calm is lost.

OpenStudy (ikram002p):

-,-

OpenStudy (kaleidoscopicsink):

Not yelled at by you. Other people.

OpenStudy (ikram002p):

ok here is away... collect all terms with c on the same side bc-dc=e-a make c a common factor c(b-d)=e-a let c alone c=(e-a)/(b-d) simple algebra !

OpenStudy (ikram002p):

now if i ask you would you do that for dx/dy you would complain and not giving a try

OpenStudy (kaleidoscopicsink):

I don't understand. I really don't.

OpenStudy (ikram002p):

ok here is another question 3x+4=x+2 solve for x

OpenStudy (kaleidoscopicsink):

I cannot do that with the given trig things. Im bad at it.

OpenStudy (ikram002p):

\(\large \begin{matrix} y \cos(x)+{\color{Red} {\dfrac{dy}{dx}}} \sin (x) =x \cos(y){\color{Red} {\dfrac{dy}{dx}}} + \sin (y) \\ {\color{Red} {\dfrac{dy}{dx}}} \sin (x)-x \cos(y){\color{Red} {\dfrac{dy}{dx}}} = \sin (y) -y \cos(x)\\ {\color{Red} {\dfrac{dy}{dx}}} [\sin (x)-x \cos(y)] = \sin (y) -y \cos(x) \\ {\color{Red} {\dfrac{dy}{dx}}} = \dfrac {\sin (y) -y \cos(x)}{ \sin (x)-x \cos(y)} \\ \end{matrix} \)

OpenStudy (kaleidoscopicsink):

Thank you so much for clarifying that!!!! My gratitude is beyond belief.

OpenStudy (ikram002p):

full solution would be like:- \( \Large \begin{matrix} (y~\sin(x))' = (x ~\sin(y) )' \\ y(\sin (x))'+y'\sin(x)=x(\sin(y))'+x'(\sin(y)) \\ y \cos(x)+\dfrac{dy}{dx} \sin (x) =x \cos(y)\dfrac{dy}{dx} + \sin (y) \\ y \cos(x)+{\color{Red} {\dfrac{dy}{dx}}} \sin (x) =x \cos(y){\color{Red} {\dfrac{dy}{dx}}} + \sin (y) \\ {\color{Red} {\dfrac{dy}{dx}}} \sin (x)-x \cos(y){\color{Red} {\dfrac{dy}{dx}}} = \sin (y) -y \cos(x)\\ {\color{Red} {\dfrac{dy}{dx}}} [\sin (x)-x \cos(y)] = \sin (y) -y \cos(x) \\ {\color{Red} {\dfrac{dy}{dx}}} = \dfrac {\sin (y) -y \cos(x)}{ \sin (x)-x \cos(y)} \\ \end{matrix} \)

OpenStudy (ikram002p):

PS:- \( \dfrac { y \cos(x)-\sin (y) }{ x \cos(y)-\sin (x)} = \dfrac {\sin (y) -y \cos(x)}{ \sin (x)-x \cos(y)} \)